# Circle packing and Riemann uniformization of random triangulations in an ergodic scale-free environment

### 06 Mar 2026 v1

@@ 1.5 Outline
In Section 4.4, we conclude our proof of Theorem 1.8 by showing that $\phi_{\infty}$ has to be the composition of a deterministic linear transform, a random rotation and a scaling. Using the polynomial growth from Lemma 4.10, the harmonic function $\phi_{\infty}\circ\varphi^{-1}$ on $\mathbb{C}$ has polynomial growth and must be a polynomial (Lemma 4.24). Furthermore, in Lemma 4.27, we prove that this polynomial must have degree 1 because if this is not the case, then one can find two cells that are close in $\mathcal{H}$ and have macroscopic distance in $\varphi(M(\mathcal{H}))$, which contradicts with the continuity as in Lemma 4.3 and Lemma 4.6. In Lemmas 4.29,  4.30 and  4.32, we further use the ergodic theorem for the cell configuration (Lemma 2.4) to show that, this linear map $\phi_{\infty}\circ\varphi^{-1}$ can be written as a composition of a deterministic linear transform $\mathbf{A}$, a random rotation and a scaling, and in Lemma 4.33, we prove that if we further assume the rotation invariance in law of $\mathcal{H}$ as in Definition 1.6, then $\mathbf{A}$ must be a rotation. This finishes the proof of Theorem 1.8. Our proof in Section 4.4 works under quite mild assumptions and gives a general method for proving that a random harmonic function is of a particular form (namely, the composition of a deterministic linear transformation and a random rotation and scaling) as long as it has polynomial growth, is almost injective, and the averages of the Dirichlet energy over large regions are approximately a deterministic constant.
In Section 5.1, we prove Theorem 5.1, a version of Theorem 1.7 and Theorem 1.8 under a weaker version of the line connectivity property. In Section 5.2, we prove a version of Theorem 1.7 for general planar maps (non-triangulations); more specifically, we consider the variant of circle packings mentioned in Section 1.4. Finally, in Section 5.3 we prove the extension of Theorem 1.8 to $p$-angulations.
@@ 2 Preliminaries
2.1 Triangulations, circle packings and Riemann surfaces
We follow the papers [BS01, AS03, GR13] for the definition on random plane triangulations. A planar map $\mathcal{M}$ is a proper (without edge-crossings) embedding of a connected planar (multi)graph $\mathcal{G}$ in the two-dimensional sphere $\widehat{\mathbb{C}}$$\mathbb{C}\cup\{\infty\}$, where we identify two planar maps if one can be obtained from the other via an orientation-preserving homeomorphism of the sphere. We will always assume that $\mathcal{M}$ is locally finite, i.e., each vertex and each face has finite degree. For a planar map $\mathcal{M}$ on the sphere $\widehat{\mathbb{C}}$$\mathbb{C}\cup\{\infty\}$, we call a connected component of $\widehat{\mathbb{C}}\backslash\mathcal{M}$ a triangular face if its boundary is consisted of precisely three edges of $\mathcal{M}$. An embedded (planar) triangulation $\mathcal{T}$ consists of a finite or infinite connected planar map $\mathcal{M}$ in the sphere, together with a subset of the triangular faces, and its support $S(\mathcal{T})\subseteq\widehat{\mathbb{C}}$ is defined to be the union of $\mathcal{M}$ and all the triangular faces in $\mathcal{M}$. Two embedded triangulations $T$ and $T^{\prime}$ are considered equivalent if there is an orientation preserving homeomorphism between their supports which maps vertices (resp., edges, faces) to vertices (resp., edges faces). For convenience we abbreviate “equivalence class of embedded triangulations” to “triangulation”. Following the terminology in [AS03, Section 1.2], we call $\mathcal{T}$ a type II triangulation if $\mathcal{T}$ has no loops but possibly multiple edges, and call $\mathcal{T}$ a type III triangulation if $\mathcal{T}$ has no loops and no multiple edges. We call a vertex $v\in\mathcal{T}$ an interior (resp. boundary) vertex of $\mathcal{T}$ if $v$ is in the interior (resp. on the boundary) of the support $S(\mathcal{T})$. We call a type II (resp. type III) triangulation $\mathcal{T}$ an infinite plane triangulation (resp. simple infinite plane triangulation), if the underlying graph is one-ended, and there are no boundary vertices. We call a type II (resp. type III) triangulation $\mathcal{T}$ a (finite) disk triangulation (resp. simple disk triangulation), if the underlying planar map $\mathcal{M}$ is finite and the support $S(\mathcal{T})$ is homeomorphic to a disk. A rooted triangulation $(\mathcal{T},e)$ is a triangulation $\mathcal{T}$ together with an oriented edge $e=(x,y)$ of $\mathcal{T}$, called the root edge. The vertex $x$ is the root vertex.
@@ 2.1 Triangulations, circle packings and Riemann surfaces
Figure 2: Two adjacent vertices $u,v$. The Dubejko conductance of the edge uvuv is defined to be $\mathfrak{c}(u,v)=\frac{|o_{a}-o_{b}|}{|o_{u}-o_{v}|}=\frac{\sqrt{r_{u}r_{v}}}{r_{u}+r_{v}}\cdot\big(\sqrt{\frac{r_{w_{1}}}{r_{w_{1}}+r_{u}+r_{v}}}+\sqrt{\frac{r_{w_{2}}}{r_{w_{2}}+r_{u}+r_{v}}}\big)$.
One can also consider random walks on circle packings. Consider the circle packing $\mathcal{P}$ of some simple infinite triangulation $\mathcal{T}$. Let $u,v$ be adjacent vertices in $\mathcal{T}$, and $w_{1},w_{2}$ be the vertices on the two faces of $\mathcal{T}$ containing the edge uvuv other than $u$ and $v$. Let $o_{u},o_{v},o_{w_{1}},o_{w_{2}}$ be the centers of the circles for $u,v,w_{1},w_{2}$, and $o_{a},o_{b}$ be the centers of circles which are inscribed to the triangles $o_{u}o_{v}o_{w_{1}}$ and $o_{u}o_{v}o_{w_{2}}$, respectively. The Dubejko conductance of the edge uvuv is defined to be $\mathfrak{c}(u,v)=\frac{|o_{a}-o_{b}|}{|o_{u}-o_{v}|}$; see Figure 2 for an illustration. If we write $r_{u},r_{v},r_{w_{1}},r_{w_{2}}$ for the radii of the disks $D_{u},D_{v},D_{w_{1}},D_{w_{2}}$, then
@@
One can infer from (2.1) that $\mathfrak{c}(u,v)<\frac{1}{2}$ for any edge uvuv using $2\sqrt{r_{u}r_{v}}\leq r_{u}+r_{v}$. Furthermore, by Dubejko’s theorem (see e.g. [STE05, Theorem 18.3]), for a circle packing $\mathcal{P}$ of a simple infinite triangulation, if we consider the random walk $(Z_{n})_{n\geq 0}$ moving between the centers of the circles in $\mathcal{P}$, then $(Z_{n})_{n\geq 0}$ is a martingale.
@@
Next we follow the description of [GR13, Section 2.2] to define the Riemann surfaces associated with triangulations. Let $\mathcal{T}$ be a disk triangulation or an infinite plane triangulation. We associate $\mathcal{T}$ a metric space $(M(\mathcal{T}),d_{M})$, by regarding each triangle of $\mathcal{T}$ as an equilateral triangle of side length 1 and defining the distance between two points as the length of the shortest path joining them. Equivalently, $M(\mathcal{T})$ is obtained by gluing equilateral triangles according to their adjacency pattern in $\mathcal{T}$. The metric space $M(\mathcal{T})$ can be equipped with a compatible Riemann surface structure by defining coordinate charts as follows. If $x$ is an interior point of a triangle or interior to an edge on the boundary of the support, then we use this triangle, viewed as a subset of the plane, as the chart, and use the identity map as the projection. Similarly, if $x$ is interior to an edge and not on the boundary of the support, then we place two adjacent equilateral triangles in the plane and use the identity as the projection. Now let $x$ be an interior vertex of degree $n$, and let $\Delta_{1},...,\Delta_{n}$ be the equilaterals which meet at $x$ in cyclic order. We place them in the plane such that for j=1,,nj=1,...,n, $\Delta_{j}$ is the triangle with vertices $0,e^{\pi(j-1)i/3},e^{\pi ji/3}$, and define the coordinate map for $z=re^{i\theta}\in\Delta_{j}$ by $\phi(z)=r^{6/n}e^{6i\theta/n}$, using $\pi(j-1)/3<\theta<\pi j/3$ on $\Delta_{j}$. By further setting $\phi(0)=0$, this defines a homeomorphism between a neighborhood of $x$ in $M(\mathcal{T})$ and a neighborhood of 0 in $\mathbb{C}$. Moreover, for the atlas chosen as above, the transition map between two overlapping charts are analytic. The case where $x$ be an boundary vertex of degree $n$ can be treated similarly using the map $z\mapsto z^{3/(n-1)}$. This defines the Riemann surface structure for $M(\mathcal{T})$. With slight abuse of language, for a vertex $v\in\mathcal{V}\mathcal{T}$, we will also view $v$ as a point on the Riemann surface $M(\mathcal{T})$.
@@
Following the uniformization theorem, for an infinite plane triangulation $\mathcal{T}$, the surface $M(\mathcal{T})$ is either conformally equivalent to the complex plane and thus parabolic, or conformally equivalent to the unit disk and thus hyperbolic. As pointed out by [GR13], an equivalent characterization is that the Brownian motion on the surface $M(\mathcal{T})$ is either recurrent (in the parabolic case) or transient (in the hyperbolic case). A function $f:M(\mathcal{T})\to\mathbb{C}$ is harmonic if and only if $f\circ\varphi^{-1}$ is a harmonic function in usual sense, where $\varphi:M(\mathcal{T})\to\mathbb{C}$ is some conformal map.

### 06 May 2026 v2

@@ 1.5 Outline
In Section 4.4, we conclude our proof of Theorem 1.7 by showing that $\phi_{\infty}$ has to be the composition of a deterministic linear transform, a random rotation and a scaling. Using the polynomial growth from Lemma 4.9, the harmonic function $\phi_{\infty}\circ\varphi^{-1}$ on $\mathbb{C}$ has polynomial growth and must be a polynomial (Lemma 4.22). Furthermore, in Lemma 4.25, we prove that this polynomial must have degree 1 because if this is not the case, then one can find two cells that are close in $\mathcal{H}$ and have macroscopic distance in $\varphi(M(\mathcal{H}))$, which contradicts with the continuity as in Lemma 4.2 and Lemma 4.5. In Lemmas 4.27,  4.28 and  4.30, we further use the ergodic theorem for the cell configuration (Lemma 2.4) to show that, this linear map $\phi_{\infty}\circ\varphi^{-1}$ can be written as a composition of a deterministic linear transform $\mathbf{A}$, a random rotation and a scaling, and in Lemma 4.31, we prove that if we further assume the rotation invariance in law of $\mathcal{H}$ as in Definition 1.5, then $\mathbf{A}$ must be a rotation. This finishes the proof of Theorem 1.7. Our proof in Section 4.4 works under quite mild assumptions and gives a general method for proving that a random harmonic function is of a particular form (namely, the composition of a deterministic linear transformation and a random rotation and scaling) as long as it has polynomial growth, is almost injective, and the averages of the Dirichlet energy over large regions are approximately a deterministic constant.
In Section 5, we extend our results to general planar maps by proving Theorem 1.8 and Theorem 1.9. In Section 5.1, we prove Theorem 1.8, where we construct a cell configuration $\widetilde{\mathcal{H}}$ in the framework of Theorem 1.6 from $\mathcal{H}$. In Section 5.2 we prove Theorem 1.9. We first add a vertex on each edge and face of the planar map $\mathcal{M}$, and further add edges to turn $\mathcal{M}$ into an infinite plane triangulation. We define a different version of semi-flowers for the vertices and the added vertices in Lemma 4.1 and prove that under conformal maps, the ratio of the outradius and inradius of the semi-flowers can be bounded by the square of the degree, as in Lemma 5.4. We also split the cell configuration $\mathcal{H}$ to get a new cell configuration whose associated map is the new infinite triangulation. This together with the same arguments in Section 4 gives the proof of Theorem 1.9. Finally, in Section 5.3, we prove Theorem 5.5, a version of Theorems 1.6-1.9 under a weaker version of the line connectivity property.
@@ 2 Preliminaries
2.1 Planar maps, circle packings and Riemann surfaces
We follow the papers [BS01, AS03, GR13] for the definition of planar maps and triangulations. A planar map $\mathcal{M}$ is a proper (in particular, without edge-crossings) embedding of a connected planar (multi)graph $\mathcal{G}$ in the two-dimensional sphere $\widehat{\mathbb{C}}$$\mathbb{C}\cup\{\infty\}$, where we identify two planar maps if one can be obtained from the other via an orientation-preserving homeomorphism of the sphere which maps vertices (resp. edges, faces) to vertices (resp. edges, faces). For convenience we abbreviate “equivalence class of embedded planar maps” to “planar map”. We will always assume that $\mathcal{M}$ is locally finite, i.e., each vertex and each face has finite degree. The face set $\mathcal{F}\mathcal{M}$ of $\mathcal{M}$ is the set of connected components of $\widehat{\mathbb{C}}\backslash\mathcal{G}$, and the support $S(\mathcal{M})\subseteq\widehat{\mathbb{C}}$ is defined to be the union of $\mathcal{M}$ and a subset of the faces in $\mathcal{F}\mathcal{M}$. In this paper we assume that all the faces in $S(\mathcal{M})$ are bounded in $\mathbb{C}$. We call a vertex $v\in\mathcal{V}\mathcal{M}$ an interior (resp. boundary) vertex of $\mathcal{M}$ if $v$ is in the interior (resp. on the boundary) of the support $S(\mathcal{M})$. We say $\mathcal{M}$ has whole-plane topology if the underlying graph is one-ended and $\mathcal{M}$ has no boundary vertices, and we say $\mathcal{M}$ has disk topology, if $\mathcal{M}$ is finite and the support $S(\mathcal{M})$ is homeomorphic to a closed disk. A rooted planar map $(\mathcal{M},e)$ is a planar map $\mathcal{M}$ together with an oriented edge $e=(x,y)$ of $\mathcal{M}$, called the root edge. The vertex $x$ is the root vertex.
For a planar map $\mathcal{M}$ on the sphere $\widehat{\mathbb{C}}$$\mathbb{C}\cup\{\infty\}$, we call a connected component of $\widehat{\mathbb{C}}\backslash\mathcal{M}$ a triangular face if its boundary is consisted of precisely three edges of $\mathcal{M}$. An embedded (planar) triangulation $\mathcal{T}$ is a finite or infinite connected planar map $\mathcal{M}$ in the sphere, such that all the faces in the support $S(\mathcal{M})$ are triangular faces. Following the terminology in [AS03, Section 1.2], we call $\mathcal{T}$ a type II triangulation if $\mathcal{T}$ has no loops but possibly multiple edges, and call $\mathcal{T}$ a type III triangulation if $\mathcal{T}$ has no loops and no multiple edges. We call a type II (resp. type III) triangulation $\mathcal{T}$ an infinite plane triangulation (resp. simple infinite plane triangulation), if $\mathcal{T}$ is a planar map with whole-plane topology. We call a type II (resp. type III) triangulation $\mathcal{T}$ a (finite) disk triangulation (resp. simple disk triangulation), if $\mathcal{T}$ is a planar map with disk topology.
@@ 2.1 Planar maps, circle packings and Riemann surfaces
Figure 2: Two adjacent vertices $u,v$. The Dubejko conductance of the edge $uv$ is defined to be $\mathfrak{c}(u,v)=\frac{|o_{a}-o_{b}|}{|o_{u}-o_{v}|}=\frac{\sqrt{r_{u}r_{v}}}{r_{u}+r_{v}}\cdot\big(\sqrt{\frac{r_{w_{1}}}{r_{w_{1}}+r_{u}+r_{v}}}+\sqrt{\frac{r_{w_{2}}}{r_{w_{2}}+r_{u}+r_{v}}}\big)$.
One can also consider random walks on circle packings. Consider the circle packing $\mathcal{P}$ of some simple infinite triangulation $\mathcal{T}$. Let $u,v$ be adjacent vertices in $\mathcal{T}$, and $w_{1},w_{2}$ be the vertices on the two faces of $\mathcal{T}$ containing the edge $uv$ other than $u$ and $v$. Let $o_{u},o_{v},o_{w_{1}},o_{w_{2}}$ be the centers of the circles for $u,v,w_{1},w_{2}$, and $o_{a},o_{b}$ be the centers of circles which are inscribed to the triangles $o_{u}o_{v}o_{w_{1}}$ and $o_{u}o_{v}o_{w_{2}}$, respectively. The Dubejko conductance of the edge $uv$ is defined to be $\mathfrak{c}(u,v)=\frac{|o_{a}-o_{b}|}{|o_{u}-o_{v}|}$; see Figure 2 for an illustration. If we write $r_{u},r_{v},r_{w_{1}},r_{w_{2}}$ for the radii of the disks $D_{u},D_{v},D_{w_{1}},D_{w_{2}}$, then
@@
One can infer from (2.1) that $\mathfrak{c}(u,v)<\frac{1}{2}$ for any edge $uv$ using $2\sqrt{r_{u}r_{v}}\leq r_{u}+r_{v}$. Furthermore, by Dubejko’s theorem (see e.g. [STE05, Theorem 18.3]), for a circle packing $\mathcal{P}$ of a simple infinite triangulation, if we consider the random walk $(Z_{n})_{n\geq 0}$ moving between the centers of the circles in $\mathcal{P}$, then $(Z_{n})_{n\geq 0}$ is a martingale.
@@
Next we define the Riemann surfaces associated with general planar maps, following the presentation in [AHN+18, Section 6]. Let $\mathcal{M}$ be a planar map with whole-plane topology. For each face $f$ of $\mathcal{M}$ of degree 1, i.e., $f$ is enclosed by a single edge of $\mathcal{M}$, we remove this edge and the face from $\mathcal{M}$. For each face $f$ of $\mathcal{M}$ of degree 2, i.e., $f$ is formed by a double edge of $\mathcal{M}$, we collapse this double edge into a single edge. Now we assume that each face of $\mathcal{M}$ has degree at least 3. We associate to $\mathcal{M}$ a metric space $(M(\mathcal{M}),d_{M})$, by regarding each face of $\mathcal{M}$ of degree $p$ as regular $p$-gon of side length 1 and defining the distance between two points as the length of the shortest path joining them. Equivalently, $M(\mathcal{M})$ is obtained by gluing regular $p$-gons of unit side length according to their adjacency pattern in $\mathcal{M}$. The metric space $M(\mathcal{M})$ can be equipped with a compatible Riemann surface structure by defining coordinate charts as follows. For a face $f$, we use the regular polygon associated with $f$, viewed as a subset of the plane, as the chart, and use the identity map as the projection. For each edge $e$ of $\mathcal{M}$, we define an open neighborhood of the interior of $e$ in $M(\mathcal{M})$ by adding to $e$ the two triangles formed by the endpoints of $e$ and the centers of the two faces adjacent to $e$. To define a coordinate chart on this neighborhood, we simply place the two triangles next to each other in the plane. The reason to take only a subset of the faces and not the entire faces is that this chart is well defined even if both sides of $e$ are incident to the same face. For each vertex $v$ of $\mathcal{M}$, we define an open neighborhood of $v$ in $M(\mathcal{M})$ similarly by intersecting the corners of the faces adjacent to $v$ with open discs of radius 1/2 centered at $v$. We define a chart on this neighborhood by first laying the corners out sequentially around the origin (with possible overlapping), and then applying the function $z\mapsto z^{2\pi/\theta(v)}$, suitably interpreted to get an injective map into the plane. The radius is chosen so that this definition remains valid when multiple corners of the same face are located at the same vertex. We will also work with the setting where $\mathcal{M}$ is a disk triangulation; in this case, the boundary vertices of degree $n$ can be treated similarly using the map $z\mapsto z^{3/(n-1)}$.
@@
Following the uniformization theorem, for a planar map $\mathcal{M}$ with whole-plane topology, the surface $M(\mathcal{M})$ is either conformally equivalent to the complex plane and thus parabolic, or conformally equivalent to the unit disk and thus hyperbolic. As pointed out by [GR13, AHN+18], an equivalent characterization is that the Brownian motion on the surface $M(\mathcal{M})$ is either recurrent (in the parabolic case) or transient (in the hyperbolic case). A function $f:M(\mathcal{M})\to\mathbb{C}$ is harmonic if and only if $f\circ\varphi^{-1}$ is a harmonic function in usual sense, where $\varphi:M(\mathcal{M})\to\mathbb{C}$ is some conformal map.